OpenStudy (anonymous):

find d^2y/dx^2 1-xy=x-y

OpenStudy (cggurumanjunath):

1-(x*y'+y)=1-y' 0-(x*y"+y'*1+y')=0-y"

OpenStudy (cggurumanjunath):

y'=dy/dx y"=d^2y/dx^2

OpenStudy (cggurumanjunath):

xy"+y'+y'=y" xy"+2y'=y" xy"-y"=-2y' y"(x-1)=-2y' y"=(-2y')/(x-1)